# ECCO: Can We Improve Model-Generated Code Efficiency Without Sacrificing Functional Correctness?

### v1

@@ 2 Related Work
It is challenging to robustly evaluate program efficiency, due to varying hardware platforms and setups. Previous works have evaluated the efficiency of code by executing code in a local software environment Singhal et al. (2024); Huang et al. (2024) or using Docker containerized environments on local hardware (Khan et al., 2023), but this can result in varying runtime and memory usage across hardware, thus causing unreliable and irreproducible efficiency evaluations. An alternative approach is to use an architecture simulator Shypula et al. (2024) which ensures the execution of each program is exactly simulated at the hardware level on the x86 architecture, but is limited to compiled languages such as C++. For more popular interpreted languages such as Python and Java (JVM), some use LeetCode’s execution engine (Niu et al., 2024; Coignion et al., 2024), but with a restricted space of testable problems. In our work, we propose an evaluation setup using an accessible cloud computing instance that ensures consistent virtual hardware and reliable benchmarking.
@@
To start, some works explore in-context learning to optimize program efficiency (Shypula et al., 2024; Huang et al., 2024). Beyond vanilla prompting, iterative prompting methods (Madaan et al., 2024; Shinn et al., 2024; Ridnik et al., 2024) have been explored to improve specific aspects of generation, by incorporating feedback or reflection from an LM or external modules. Meanwhile, Shypula et al. (2024) propose fine-tuning with self-play synthetic data, with retrieved in-context examples via a nearest neighbor search. However, none of these methods have been rigorously studied in the correctness-preserving optimization setup. We fill in this gap and provide systematic studies of all methods.
@@ 3.1 Evaluation Platform
Figure 2: Our evaluation platform using Judge0.
@@ 3.3.2 NL-Instructed Generation
Figure 4: Illustration of NL-based generation.
@@ 4.1 In-Context Learning
Many LMs perform better when incorporating instructions Ouyang et al. (2022); Wei et al. (2022). We use two prompts: $I_{gen}$ for NL-based generation which instructs models to generate correct and efficient programs; and $I_{eff}$ for history-based editing which instructs models to optimize the input program. $I_{eff}$ is adapted from PIE Shypula et al. (2024) and Self-Refine Madaan et al. (2024). See §C for details.
@@ 4.2 Iterative Refinement
We explore three methods to iteratively improve and refine the generated code to be more efficient, which intuitively aligns with the way that humans continuously refine code (Madaan et al., 2024).
@@
4.3 Fine-tuning
Vanilla Fine-tuning
Execution Conditioned Fine-tuning
Trajectory Conditioned Fine-tuning
@@ 4.3 Fine-tuning
Table 2: Generation results using in-context learning approaches (instruction-prompting and few-shot examples).
@@ 5.1 Experimental Setup
We experiment with several best-performing LMs pre-trained on code. Specifically, we evaluate CodeLlama-13B Roziere et al. (2023), DeepSeekCoder-7B Guo et al. (2024), CodeGemma-7B Team et al. (2024), WizardCoder-13B-Python Luo et al. (2023), StarCoder2-15B Lozhkov et al. (2024). We use the instruction-tuned versions of all of these open-checkpoint models unless indicated otherwise. We also use the proprietary GPT-4o model for no-training methods.
@@ 5.2.1 In-Context Learning
As shown in Table 2, all methods either reduce pass@1 by a large margin (in editing mode) or obtain low pass@1 (generation mode). Comparing the two paradigms, history-based editing results in a substantially higher pass@1 by referring to a base correct program, compared to NL-instructed generation which lacks a base program to start from. GPT-4o obtains a much higher pass@1 than all models in both paradigms.
@@
Table 3: Generation results with iterative refinement approaches for two optimization paradigms.
@@ 6 Additional Analysis
Multiple refinement iterations may improve results by allowing more turns for models to refine. To verify this, we evaluate the iterative refinement method using 1–4 iterations. While self-refine and exec-refine improve program speedup over iterations (5(b)), all methods continuously degrade the pass@1 to various extents. For memory usage (5(c)), exec-refine consistently reduces memory usage, yet other methods exhibit big fluctuations. In general, more iterations can speed up the program, yet further sacrifice correctness, thus has limited gains in terms of correctness-preserving optimization.
@@
To study whether models can recover from incorrect starting solutions in the history-based editing paradigm, we evaluate on a collection of 157 pairs of programs, ECCO-Fix, where the input code is almost correct: one that passed all public test cases but fails a few private test cases. As shown by Table 7, we show that exec-refine can fix incorrect programs to pass all public and private test cases, with access to only the PASS/FAIL status of public test cases. In comparison, self-refine breaks the correctness of more programs. Aligning with our findings in earlier sections and §A, exec-refine, with execution information in contexts, can encourage models to generate functionally correct programs.
@@ Acknowledgements
We thank Aman Madaan for the helpful discussions.
@@ Appendix A Iterative Refinement Analysis
Table 8: Comparing iterative refinement approaches on % Optimization metrics. Exec-Refine is the best performing approach across all models and metrics
@@ Appendix B Implementation Details
Appendix C Prompt Details
We illustrate and detail all of the prompts used for the experiments below.
@@ Appendix C Prompt Details
Write a python solution for the problem below which is
correct and efficient (runtime and memory).
@@
{self-feedback / execution-feedback / reflection}

### v2

@@ 2 Related Work
It is challenging to robustly evaluate program efficiency, due to varying hardware platforms and setups. Previous works have evaluated the efficiency of code by executing code in a local software environment Singhal et al. (2024); Huang et al. (2024) or using containerized environments on local hardware (Khan et al., 2023), but this can result in varying runtime and memory usage across hardware, thus causing irreproducible evaluations. An alternative approach is to use an architecture simulator Shypula et al. (2024) which ensures the execution of each program is exactly simulated at the hardware level, but is limited to compiled languages such as C++. A hardware counter though reliable Liu et al. (2024), cannot measure memory usage. For popular interpreted languages such as Python and Java, some use LeetCode’s execution engine (Niu et al., 2024; Coignion et al., 2024), but with a restricted space of testable problems. In our work, we propose an evaluation setup using an accessible cloud computing instance that ensures consistent virtual hardware and reliable benchmarking.
@@
To start, some works explore in-context learning to optimize program efficiency (Huang et al., 2024), with retrieval methods to select relevant examples (Gao et al., 2024; Shypula et al., 2024). Beyond vanilla prompting, iterative prompting methods (Madaan et al., 2024; Shinn et al., 2024; Ridnik et al., 2024) have been explored to improve specific aspects of generation, by incorporating feedback from an LM or external modules. Meanwhile, finetuning has been proposed with self-play, synthetic preferences and problem-oriented data Shypula et al. (2024); Gee et al. (2024); Ye et al. (2024). However, none of these methods have been rigorously studied for correctness-preserving optimization. We fill in this gap and provide systematic studies of all methods.
@@ 3.1 Evaluation Platform
Figure 2: Evaluation platform using Judge0.
@@ 3.3.2 NL-Instructed Generation
Figure 4: Illustration of NL-instructed generation.
@@ 3.4 ECCO Feature Analysis
Figure 5: Iterative refinement methods utilizing different forms of feedback. Self-Refine uses Natural Language feedback, Exec-Refine uses raw execution results on $T_{public}$ and NL+Exec-Refine uses NL reflection of execution.
@@ 4.1 In-Context Learning
Many LMs perform better when incorporating instructions Ouyang et al. (2022); Wei et al. (2022). We use two prompts: $I_{gen}$ for NL-based generation which instructs models to generate correct and efficient programs; and $I_{eff}$ for history-based editing which instructs models to optimize the input program. $I_{eff}$ is adapted from PIE Shypula et al. (2024) and Self-Refine Madaan et al. (2024). See §D for details.
@@ 4.2 Iterative Refinement
We explore three methods (illustrated in Figure 5 ) to iteratively refine the generated code to be more efficient, which intuitively aligns with the way that humans improve code (Madaan et al., 2024).
@@
Table 2: Results using In-Context Learning approaches (instruction-prompting and few-shot learning)
4.3 Fine-tuning
Vanilla Fine-tuning
Execution Conditioned Fine-tuning
Trajectory Conditioned Fine-tuning
@@ 5.1 Experimental Setup
We experiment with several best-performing LMs pre-trained on code. Specifically, we evaluate CodeLlama-13B Roziere et al. (2023), DeepSeekCoder-7B-v1.5 Guo et al. (2024), CodeGemma-7B Team et al. (2024), WizardCoder-13B-Python Luo et al. (2023), StarCoder2-15B Lozhkov et al. (2024). We use the instruction-tuned versions of all of these open-checkpoint models unless indicated otherwise. We also use the proprietary GPT-4o model for no-training methods.
@@ 5.2.1 In-Context Learning
As shown in Table 2, all methods either reduce pass@1 of the program by a large margin (in editing mode) or obtain low pass@1 (generation mode). Comparing the two paradigms, history-based editing results in a substantially higher pass@1 by referring to a base correct program, compared to NL-instructed generation which lacks a base program to start from. GPT-4o obtains a much higher pass@1 than all models in both paradigms.
@@
Table 3: Results with iterative refinement approaches. Feedback in NL (self & nl+exec) improves efficiency better, whereas raw execution feedback (exec) maintains correctness more effectively.
@@ 6 Additional Analysis
Multiple refinement iterations may improve results by allowing more turns for models to refine. To verify this, we evaluate the iterative refinement method using 1–4 iterations. While self-refine and exec-refine improve program speedup over iterations (6(b)), all methods continuously degrade the pass@1 to various extents. Exec-refine preserves correctness more effectively in further iterations as well. For memory usage (6(c)), exec-refine consistently reduces memory usage, yet other methods exhibit big fluctuations. In general, more iterations can speed up the program, yet further sacrifice correctness, thus has limited gains in terms of correctness-preserving optimization.
Figure 7: Correctness and efficiency of generated code across model sizes for CodeLLaMa and DeepseekCoder. The top row corresponds to History-Based Editing and the bottom row includes NL-Instructed Generation.
@@
To study whether models can recover from incorrect starting solutions in the history-based editing paradigm, we evaluate on a collection of 157 pairs of programs, ECCO-Fix, where the input code is almost correct: one that passed all public test cases but fails a few private test cases. As shown by Table 7, we show that exec-refine can fix incorrect programs to pass all public and private test cases, with access to only the PASS/FAIL status of public test cases. In comparison, self-refine breaks the correctness of more programs.
Aligning with our findings in earlier sections and §A, exec-refine, with execution information in contexts, can encourage models to generate functionally correct programs.
@@
How does scale of model impact efficiency and correctness?
We perform experiments across model scales for the CodeLLaMa (7B, 13B, 34B, 70B) and DeepseekCoder-v1 (1.3B, 6.7B, 33B) instruction-tuned model families. As seen in Figure 7 and Table 10, for both History-based Editing and NL-Instructed Generation, scaling model size presents a mix of benefits and trade-offs in terms of efficiency and correctness. For History-based Editing, larger CodeLLaMa models (such as 70B) show better speedup but experience diminishing gains in correctness (Pass@1), with a slight decrease in memory reduction. DeepseekCoder follows a similar pattern but consistently underperforms CodeLLaMa in correctness. In NL-Instructed Generation, scaling models significantly improves correctness, as seen with CodeLLaMa 70B and DeepseekCoder 33B, although giving varying results for runtime and memory percentiles.
@@ Acknowledgements
We thank Aman Madaan for the helpful discussions in the early stage of the project.
@@ Appendix A Iterative Refinement Analysis
Table 8: Comparing iterative refinement approaches on % Optimization metrics. Exec-Refine is the best performing approach across all models and metrics.
@@ Appendix B Implementation Details
Appendix C Variance Analysis
We measured the variance for two models (DeepseekCoder and CodeLLaMa) by sampling generations three times and evaluating each generation independently on our Judge setup. The results are included below with 95% Confidence Intervals are shown in Table 9.
Pass@1 has no variance across runs in both paradigms due to the low temperature; runtime efficiency metrics (speedup and runtime %) both only have about  1% Confidence Interval (CI). Memory has close to no variance in the history-based editing setting but shows  5% CI in NL-instructed generation.
Model History-based Editing NL-Instructed Generation
Pass@1 Speedup Memory Reduction Pass@1 Runtime (%) Memory (%)
DeepseekCoder 32.87 $\pm$ 0.00 2.112 $\pm$ 0.022 1.097 $\pm$ 0.000166 18.75 $\pm$ 0.00 57.589 $\pm$ 1.012 67.796 $\pm$ 3.106
CodeLLaMa 57.50 $\pm$ 0.00 1.456 $\pm$ 0.0028 1.114 $\pm$ 0.00173 8.30 $\pm$ 0.00 55.663 $\pm$ 0.720 61.029 $\pm$ 3.818
Table 9: Variance of results for History-based Editing and NL-Instructed Generation
Model Size History-based Editing NL-instructed Generation
Pass@1 Speedup Memory Reduction Pass@1 Runtime (%) Memory (%)
CodeLLaMa 7B 58.4 1.04 1.00   2.1 39.73 26.93
13B 57.5 1.44 1.11   8.3 45.30 74.18
34B 51.9 1.31 1.13   8.3 43.23 70.16
70B 55.3 1.73 1.09 12.5 39.15 64.89
DeepseekCoder 1.3B 44.2 1.36 1.13   8.3 46.71 65.62
6.7B 34.8 1.47 1.16 29.2 33.48 76.19
33B 49.1 1.51 1.09 37.5 43.91 75.61
Table 10: Results for History-based Editing and NL-Instructed Generation across different model sizes.
Appendix D Prompt Details
We illustrate and detail all of the prompts used for the experiments in Figures 8-14.
@@ Appendix D Prompt Details
Write a python code which is correct and efficient in terms of runtime and memory
usage for the following problem description.
@@
Give feedback in english for why the code solution below is incorrect
or inefficient and how the program can be fixed.
## Candidate solution:
{most recent code attempt}
## Feedback for incorrectness/inefficiency and how it can be improved:
Figure 11: Prompt used for NL-reasoning to get feedback
{mdframed}
Refine the given incorrect or sub-optimal code solution based on the feedback
specified below.
@@
{self-feedback / execution-feedback / NL+Exec-Refine}